What is the perimeter of an equilateral triangle whose one side is 5x-2 and another side that is equal to x
Answer by srinivas.g(540)

In Equilateral triangle, all sides are equal
hence 5x-2 =x
move x to the left
5x-2-x=0
move -2 to the right
5x-x= 2
4x= 2
divide with 2 on both sides
4x/4 =2/4
x= 1/2
Perimeter =sum of all sides
= x+x+x ( as all sides are equal)
=3x
= 3*(1/2)
=3/2
Perimeter = 3/2